OpenStudy (anonymous):

a store mixes Brazilian coffee worth $12 per kg and Venezuelan coffee worth $15 per kg. the mixture is to sell for $14 per kg. Find how much each should be used to make 348 kg mixture.

OpenStudy (anonymous):

First, use a weighted average to find the relative amounts of Brazilian (b) and Venezuelan (v) coffee. They will be percents, so b+v=1. 12b + 15v = 14 Once you have those percents, you can use them to find out how much of each is in the 348kg mixture.

OpenStudy (anonymous):

x=116 y=232?

OpenStudy (anonymous):

Which one is B and which is V?

OpenStudy (anonymous):

i think v=116 b=232? can someone verify with me

OpenStudy (anonymous):

Well, telling you that those aren't right will kinda' tell you what is right. But, no, those are backwards. Can you see why?

OpenStudy (anonymous):

oh yeah my bad I wrote it down the right way. just didnt type it the right way

OpenStudy (anonymous):

I'm going to try a slight variation on dmancine's method; let's see... Let b=pounds of Brazilian coffee needed let v==pounds of Venzuaelan coffee needed then the two can be related by total weight: b+v=348 [eqn 1] they can also be reltaed by cost 12b+15v=14(b+v) or -2b+v=0 solving for v (to use substitution) v=2b [eqn 2] Now sub 2b for v in [eqn 1] b+2b=348 3b=348 b=116 Now sub b=116 into [eqn 2] to find v v=2(116)=232 So, b=116, and v=232

OpenStudy (anonymous):

Here's a simple picture to see what's going on with the weighted average. The desired mix is 2/3 of the way from B to V, so you need more V to pull it closer. The relative distances tell you the relative amounts of each to include in the mix (1/3 and 2/3), and which one it's closer to tells you which one needs to be bigger (closer to V, so v is 2/3).